OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2).

OpenStudy (anonymous):

paralel, means same slope: 3x − y = −3 reorganizing: y=3x+3 slope is 3 so: y+2=3(x-4) y=3x-14